# Scalable Composite Likelihood Estimation of Probit Models with Crossed Random Effects

### v2

@@ 3.2 Robust sandwich variance
After a customary Taylor approximation, the variance of $\hat{\bm{\theta}}$ is $\mathrm{var}(\hat{\bm{\theta}})\doteq D\,\mathrm{var}(\hat{\bm{\psi}})\,D^{\mathsf{T}},$ with $D$ the Jacobian matrix of the reparameterization from $\bm{\theta}$ to $\bm{\psi}$. Letting $\bm{u}_{\text{arc}}(\bm{\psi})$ denote the score vector of the ARC estimator constructed by stacking the scores of the three misspecified likelihoods, the asymptotic variance of $\hat{\bm{\psi}}$ is
$\mathrm{avar}(\hat{\bm{\psi}})=\mathcal{I}_{\text{arc}}^{-1}(\bm{\psi}_{0})V_{\text{arc}}(\bm{\psi}_{0})\mathcal{I}_{\text{arc}}^{-1}(\bm{\psi}_{0}),$ (12)
@@
where $\mathcal{I}_{\text{arc}}({\bm{\psi}})=-\mathrm{E}(\partial\bm{u}_{\text{arc}}({\bm{\psi}})/\partial{\bm{\psi}})$ and $V_{\text{arc}}({\bm{\psi}})=\mathrm{var}(\bm{u}_{\text{arc}}({\bm{\psi}}))$ are the expected information and the score variance for the ARC estimator. These two matrices are not equal because the second Bartlett identity does not hold for misspecified likelihoods that constitute the ARC method. While estimation of the ‘bread’ matrix $\mathcal{I}_{\text{arc}}$ of the sandwich is not problematic, direct computation of the ‘filling‘ matrix $V_{\text{arc}}$ is not feasible in our large-scale setup because it requires us to approximate a large number of multidimensional integrals with a cost that does not meet our $\mathcal{O}(N)$ constraint.
@@
Since ARC estimates require $\mathcal{O}(N)$ computations, we can estimate the variance of $\hat{\bm{\theta}}$ with a parametric bootstrap. However, it is preferable to evaluate the estimation uncertainty without assuming the correctness of the fitted model, for example using the nonparametric pigeonhole bootstrap described in Owen (2007). In that approach, the rows in the data set are resampled independently of the columns. So if a row is included twice and a column is included three times, the corresponding element is included six times. The resulting bootstrap variance for a mean (such as one in a score equation) over-estimates the random effects variance by an asymptotically negligible amount. It does not require homoscedasticity of either the random effects or the errors.
@@
Now we describe a convenient approach that we have developed to estimate the variance of $\hat{\bm{\beta}}$ in $\mathcal{O}(N)$ operations without the need of resampling and repeated fitting as in the bootstraps mentioned above. The partitioned expected ARC information matrix is
@@
$\mathcal{I}_{\text{arc}}(\bm{\psi})$ $=-\begin{pmatrix}\mathrm{E}(\partial^{2}\ell_{\text{all}}/\partial\bm{\gamma}\partial\bm{\gamma}^{\mathsf{T}})&0&0\\ \mathrm{E}(\partial^{2}\ell_{\text{row}}/\partial\bm{\gamma}\partial\tau^{2}_{A})&\mathrm{E}(\partial^{2}\ell_{\text{row}}/\partial\tau^{2}_{A}\partial\tau^{2}_{A})&0\\ \mathrm{E}(\partial^{2}\ell_{\text{col}}/\partial\bm{\gamma}\partial\tau^{2}_{B})&0&\mathrm{E}(\partial^{2}\ell_{\text{col}}/\partial\tau^{2}_{B}\partial\tau^{2}_{B})\\ \end{pmatrix}.$
@@
Since the expected ARC information is triangular, then the asymptotic variance for $\hat{\bm{\gamma}}$ is
$\mathrm{avar}(\hat{\bm{\gamma}})=\mathcal{I}_{\text{all}}^{-1}(\bm{\gamma}_{0})V_{\text{all}}(\bm{\gamma}_{0})\mathcal{I}_{\text{all}}^{-1}(\bm{\gamma}_{0}),$ (13)
@@
where $\mathcal{I}_{\text{all}}({\bm{\gamma}})=-\mathrm{E}(\partial^{2}\ell_{\text{all}}/\partial\bm{\gamma}\partial\bm{\gamma}^{\mathsf{T}})$ and $V_{\text{all}}({\bm{\gamma}})=\mathrm{var}(\partial\ell_{\text{all}}/\partial\bm{\gamma})$ are the Fisher expected information and the score variance of the marginal probit model. The asymptotic variance (13) for $\hat{\bm{\gamma}}$ is thus the same as that of the estimator that maximizes the all likelihood when the nuisance parameters $\tau^{2}_{A}$ and $\tau^{2}_{B}$ are known. The robust sandwich estimator of the variance of $\hat{\bm{\gamma}}$ is obtained by replacing $\mathcal{I}_{\text{all}}(\bm{\gamma}_{0})$ and $V_{\text{all}}(\bm{\gamma}_{0})$ with some estimators that are consistent and robust to misspecification. The expected information is naturally estimated with the observed information,
$J_{\text{all}}(\hat{\bm{\gamma}})=\sum_{ij}Z_{ij}\varphi(\hat{\eta}_{ij})\left(y_{ij}\frac{\varphi(\hat{\eta}_{ij})+\hat{\eta}_{ij}\,\Phi(\hat{\eta}_{ij})}{\Phi(\hat{\eta}_{ij})^{2}}+(1-y_{ij})\frac{\varphi(\hat{\eta}_{ij})-\hat{\eta}_{ij}\,\Phi(-\hat{\eta}_{ij})}{\Phi(-\hat{\eta}_{ij})^{2}}\right)\bm{x}_{ij}\bm{x}_{ij}^{\mathsf{T}},$
@@
where $\hat{\eta}_{ij}=\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}.$ Estimation of $V_{\text{all}}(\bm{\gamma}_{0})$ is more involved. This matrix can be decomposed into the sum of three terms,
$V_{\text{all}}(\bm{\gamma}_{0})$ $=\sum_{ij}\sum_{rs}Z_{ij}Z_{rs}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{rs}^{\mathsf{T}}(\bm{\gamma}_{0})\right)$
$=\sum_{ijs}Z_{ij}Z_{is}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{is}^{\mathsf{T}}(\bm{\gamma}_{0})\right)+\sum_{ijr}Z_{ij}Z_{rj}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{rj}^{\mathsf{T}}(\bm{\gamma}_{0})\right)$
$-\sum_{ij}Z_{ij}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{ij}^{\mathsf{T}}(\bm{\gamma}_{0})\right),$
where $\bm{u}_{ij}(\bm{\gamma})$ is the score for a single observation $Y_{ij},$
$\bm{u}_{ij}(\bm{\gamma})=\frac{\varphi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})\left(y_{ij}-\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})\right)\bm{x}_{ij}}{\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})\Phi(-\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})}.$
The corresponding estimator of $V_{\text{all}}(\bm{\gamma}_{0})$ is
$\widehat{V}_{\text{all}}(\hat{\bm{\gamma}})=\widehat{V}_{A}(\hat{\bm{\gamma}})+\widehat{V}_{B}(\hat{\bm{\gamma}})-\widehat{V}_{A\cap B}(\hat{\bm{\gamma}}),$
@@
$\widehat{V}_{A}(\hat{\bm{\gamma}})=\sum_{i}\bm{u}_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}})\bm{u}_{i\text{\tiny$\bullet$}}^{\mathsf{T}}(\hat{\bm{\gamma}}),\quad\widehat{V}_{B}(\hat{\bm{\gamma}})=\sum_{j}\bm{u}_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}})\bm{u}_{\text{\tiny$\bullet$}j}^{\mathsf{T}}(\hat{\bm{\gamma}}),\quad\widehat{V}_{A\cap B}(\hat{\bm{\gamma}})=\sum_{ij}Z_{ij}\bm{u}_{ij}(\hat{\bm{\gamma}})\bm{u}_{ij}^{\mathsf{T}}(\hat{\bm{\gamma}}),$
where $\bm{u}_{i\text{\tiny$\bullet$}}(\bm{\gamma})=\sum_{j}Z_{ij}\bm{u}_{ij}(\bm{\gamma})$ and $\bm{u}_{\text{\tiny$\bullet$}j}(\bm{\gamma})=\sum_{i}Z_{ij}\bm{u}_{ij}(\bm{\gamma}).$ Estimators of the form $\widehat{V}_{\text{all}}(\hat{\bm{\gamma}})$ are used in statistical modeling of data clustered within multiple levels in medical applications (Miglioretti and Heagerty, 2004) and in economics (Cameron et al., 2011) where they are known as two-way cluster-robust sandwich estimators.
Finally, we approximate the variance of $\hat{\bm{\beta}}$ by plugging in the estimates $\widehat{\sigma}_{A}^{2}$ and $\widehat{\sigma}_{B}^{2}$,
$\widehat{\mathrm{var}}(\hat{\bm{\beta}})=(1+\widehat{\sigma}_{A}^{2}+\widehat{\sigma}_{B}^{2})\,\widehat{\mathrm{var}}(\hat{\bm{\gamma}})=(1+\widehat{\sigma}_{A}^{2}+\widehat{\sigma}_{B}^{2})\,{J}_{\text{all}}^{-1}(\hat{\bm{\gamma}})\widehat{V}_{\text{all}}(\hat{\bm{\gamma}}){J}_{\text{all}}^{-1}(\hat{\bm{\gamma}}).$ (14)
@@ 4 Simulations
In this section, we simulate from the probit model with crossed random effects (1) and compare the performance of the ARC estimator with the traditional estimator obtained by maximizing the first-order Laplace approximation of the likelihood. A further method considered is given by an infeasible oracle estimator that uses the unknown true values of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ to estimate the regression parameters as $\hat{\bm{\beta}}_{\mathrm{oracle}}=\hat{\bm{\gamma}}(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}.$ The ARC method instead corrects $\hat{\bm{\gamma}}$ using estimates of the variance components. All methods were implemented in the R language. The package TMB (Kristensen et al., 2016) was used for the Laplace approximation, with the nlminb optimization function employed for its maximization. We did not use the popular glmer function from the R package lme4 (Bates et al., 2015) because TMB is substantially more computationally efficient and thus allowed us to compare our method to the first-order Laplace approximation at larger dimensions than it would otherwise be possible with glmer. The row- and column-wise likelihoods were coded in C++ and integrated in R with Rcpp (Eddelbuettel, 2013), and optimized by Brent’s method as implemented in the optimise function. The code is available in the public repository https://github.com/rugbel/arcProbit.
@@ 4.1 Simulation settings
We considered eight different settings defined by combining three binary factors. The first factor is whether the simulation is balanced (equal numbers of rows and columns) or imbalanced (with very unequal numbers) like we typically see in applications. The second factor is whether the regression model is null apart from a nonzero intercept or has nonzero regression coefficients. The third factor is whether the random effect variances are set at a high level or at a low level. Given $R$ and $C$, the set $\mathcal{S}$ is obtained by IID Bernoulli sampling with probability $N/(RC)$. This makes the attained value of $N$ random but with a very small coefficient of variation.
We denote by $R=N^{\rho}$ and $C=N^{\kappa}$ the number of rows and columns in the data, as a power of the total sample size $N$. The two levels of the balance factor are
 Balanced
$\rho=\kappa=0.56$, and
 Imbalanced
$\rho=0.88$ and $\kappa=0.53$.
@@
This balanced case was already considered in Ghosh et al. (2022b). Because $\rho+\kappa>1$, the fraction of possible observations in the data is $N/(RC)=N^{1-\rho-\kappa}\to 0$ as $N\to\infty$ providing asymptotic sparsity in both cases. While the first choice has $R/C$ constant, the second choice has $R/C\to\infty$ with $N$. We believe that this asymptote is a better description of our motivating problems than either a setting with $C$ fixed as $R\to\infty$ or the setting common in random matrix theory (Edelman and Rao, 2005) where $R$ and $C$ diverge with $R/C$ approaching a constant value.
@@
We considered seven predictors generated from a multivariate zero-mean normal distribution with a covariance matrix $\Sigma$ corresponding to an autocorrelation process of order one, so that the entry $(k,l)$ of $\Sigma$ is $\phi^{|k-l|}$. We set $\phi=0.5$ in all the simulations. We always used the intercept $\beta_{0}=-1.2$ because in our applications $\Pr(Y=1)<1/2$ is typical. For the predictor coefficients we considered the following two choices
 Null
$\beta_{\ell}=0$ for $\ell=1,\ldots 7$, and
 Linear
$\beta_{\ell}=-1.2+0.3\ell$ for $\ell=1,\dots,7.$
The first setting is a null one where $\bm{x}$ is not predictive at all, while the second setting has modestly important nonzero predictors whose values are in linear progression.
The two choices for the variance component parameters are
 High variance
$\sigma_{A}=1$ and $\sigma_{B}=1$, and
 Low variance
$\sigma_{A}=0.5$ and $\sigma_{B}=0.2$.
@@
We chose the first setting to include variances higher than what is typically observed in applications. The second setting is closer to what we have seen in data such as that in Section 5. We represent the eight settings with mnemonics as shown in Table 1. For example, Imb-Nul-Hi  means row-column imbalance ($\rho=0.88,\kappa=0.53$), all predictor coefficients are zero and the main effect variances are large ($\sigma_{A}=1,\sigma_{B}=1$).

### v4

@@ 2.3 Robust sandwich variance
After a customary Taylor approximation, the variance of $\hat{{\theta}}$ is $\mathrm{var}(\hat{{\theta}})\doteq D\,\mathrm{var}(\hat{{\psi}})\,D^{\mathrm{\scriptscriptstyle T}},$ with $D$ the Jacobian matrix of the reparameterization from ${\theta}$ to ${\psi}$. Letting $\psi_{0}$ denote the true value of $\psi$ and ${u}_{\text{arc}}({\psi})$ denote the score vector of the all-row-column estimator constructed by stacking the scores of the three misspecified likelihoods, the asymptotic variance of $\hat{{\psi}}$ is
$\mathrm{avar}(\hat{{\psi}})=\mathcal{I}_{\text{arc}}^{-1}({\psi}_{0})V_{\text{arc}}({\psi}_{0})\mathcal{I}_{\text{arc}}^{-1}({\psi}_{0}),$ (11)
@@
where $\mathcal{I}_{\text{arc}}(\psi)=-\mathrm{E}\{\partial{u}_{\text{arc}}(\psi)/\partial\psi\}$ and $V_{\text{arc}}(\psi)=\mathrm{var}\{u_{\text{arc}}(\psi)\}$ are the expected information and the score variance for the all-row-column estimator. These two matrices are not equal because the second Bartlett identity does not hold for the misspecified likelihoods that constitute the all-row-column method. While estimation of the ‘bread’ matrix $\mathcal{I}_{\text{arc}}$ of the sandwich is not problematic, direct computation of the ‘filling’ matrix $V_{\text{arc}}$ is not feasible in our large-scale setup because it requires us to approximate a large number of multidimensional integrals with a cost that does not meet our $O(N)$ constraint.
@@
Since all-row-column estimates require $O(N)$ computations, we can estimate the variance of $\hat{\theta}$ with a parametric bootstrap. However, it is preferable to evaluate the estimation uncertainty without assuming the correctness of the fitted model, for example using the nonparametric pigeonhole bootstrap described in Owen (2007). In that approach, the rows in the data set are resampled independently of the columns. So if a row is included twice and a column is included three times, the corresponding element is included six times. The resulting bootstrap variance for a mean (such as one in a score equation) over-estimates the random effects variance by an asymptotically negligible amount. It does not require homoscedasticity of either the random effects or the errors.
@@
Now we describe a convenient approach that we have developed to estimate the variance of $\hat{{\beta}}$ in $O(N)$ operations without the need of resampling and repeated fitting as in the bootstraps mentioned above. The partitioned expected all-row-column information matrix is
@@
$\mathcal{I}_{\text{arc}}({\psi})$ $=-\begin{pmatrix}\mathrm{E}(\partial^{2}\ell_{\text{all}}/\partial{\gamma}\partial{\gamma}^{\mathrm{\scriptscriptstyle T}})&0&0\\[5.0pt] \mathrm{E}(\partial^{2}\ell_{\text{row}}/\partial{\gamma}\partial\tau^{2}_{A})&\mathrm{E}(\partial^{2}\ell_{\text{row}}/\partial\tau^{2}_{A}\partial\tau^{2}_{A})&0\\[5.0pt] \mathrm{E}(\partial^{2}\ell_{\text{col}}/\partial{\gamma}\partial\tau^{2}_{B})&0&\mathrm{E}(\partial^{2}\ell_{\text{col}}/\partial\tau^{2}_{B}\partial\tau^{2}_{B})\\ \end{pmatrix},$
@@
where $\ell_{\text{all}}=\log L_{\text{all}}(\gamma)$ from (5), $\ell_{\text{row}}=\log L_{\text{row}}(\tau^{2}_{A})$ from (9) and $\ell_{\text{col}}=\log L_{\text{col}}(\tau^{2}_{B})$. Since $\mathcal{I}_{\text{arc}}({\psi})$ is triangular, then the asymptotic variance for $\hat{{\gamma}}$ is
$\mathrm{avar}(\hat{{\gamma}})=\mathcal{I}_{\text{all}}^{-1}({\gamma}_{0})V_{\text{all}}({\gamma}_{0})\mathcal{I}_{\text{all}}^{-1}({\gamma}_{0}),$ (12)
@@
where $\mathcal{I}_{\text{all}}({{\gamma}})=-\mathrm{E}(\partial^{2}\ell_{\text{all}}/\partial{\gamma}\partial{\gamma}^{\mathrm{\scriptscriptstyle T}})$ and $V_{\text{all}}({{\gamma}})=\mathrm{var}(\partial\ell_{\text{all}}/\partial{\gamma})$ are the Fisher expected information and the score variance of the all likelihood. The asymptotic variance (12) for $\hat{{\gamma}}$ is thus the same as that of the estimator that maximizes the all likelihood when the nuisance parameters $\tau^{2}_{A}$ and $\tau^{2}_{B}$ are known. The robust sandwich estimator of the variance of $\hat{{\gamma}}$ is obtained by replacing $\mathcal{I}_{\text{all}}({\gamma}_{0})$ and $V_{\text{all}}({\gamma}_{0})$ with some estimators that are consistent and robust to misspecification. The expected information is naturally estimated with the observed information,
$J_{\text{all}}(\hat{{\gamma}})=\sum_{ij}Z_{ij}\varphi(\hat{\eta}_{ij})\biggl{\{}y_{ij}\frac{\varphi(\hat{\eta}_{ij})+\hat{\eta}_{ij}\,\Phi(\hat{\eta}_{ij})}{\Phi(\hat{\eta}_{ij})^{2}}+(1-y_{ij})\frac{\varphi(\hat{\eta}_{ij})-\hat{\eta}_{ij}\,\Phi(-\hat{\eta}_{ij})}{\Phi(-\hat{\eta}_{ij})^{2}}\biggr{\}}{x}_{ij}{x}_{ij}^{\mathrm{\scriptscriptstyle T}},$
@@
where $\hat{\eta}_{ij}={x}_{ij}^{\mathrm{\scriptscriptstyle T}}\hat{{\gamma}}.$ Estimation of $V_{\text{all}}({\gamma}_{0})$ is more involved. This matrix can be decomposed into the sum of three terms,
$V_{\text{all}}({\gamma}_{0})$ $=\sum_{ij}\sum_{rs}Z_{ij}Z_{rs}\mathrm{E}\bigl{\{}{u}_{ij}({\gamma}_{0}){u}_{rs}^{\mathrm{\scriptscriptstyle T}}({\gamma}_{0})\bigr{\}}$
$=\sum_{ijs}Z_{ij}Z_{is}\mathrm{E}\bigl{\{}{u}_{ij}({\gamma}_{0}){u}_{is}^{\mathrm{\scriptscriptstyle T}}({\gamma}_{0})\bigr{\}}+\sum_{ijr}Z_{ij}Z_{rj}\mathrm{E}\bigl{\{}{u}_{ij}({\gamma}_{0}){u}_{rj}^{\mathrm{\scriptscriptstyle T}}({\gamma}_{0})\bigr{\}}$
$-\sum_{ij}Z_{ij}\mathrm{E}\bigl{\{}{u}_{ij}({\gamma}_{0}){u}_{ij}^{\mathrm{\scriptscriptstyle T}}({\gamma}_{0})\bigr{\}},$
where ${u}_{ij}({\gamma})$ is the score for a single observation $Y_{ij},$
${u}_{ij}({\gamma})=\frac{\varphi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})\bigl{\{}y_{ij}-\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})\bigr{\}}{x}_{ij}}{\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})\Phi(-{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})}.$
The corresponding estimator of $V_{\text{all}}({\gamma}_{0})$ is
$\hat{V}_{\text{all}}(\hat{{\gamma}})=\hat{V}_{A}(\hat{{\gamma}})+\hat{V}_{B}(\hat{{\gamma}})-\hat{V}_{A\cap B}(\hat{{\gamma}}),$
@@
$\hat{V}_{A}(\hat{{\gamma}})=\sum_{i}{u}_{i\text{\tiny$\bullet$}}(\hat{{\gamma}}){u}_{i\text{\tiny$\bullet$}}^{\mathrm{\scriptscriptstyle T}}(\hat{{\gamma}}),\quad\hat{V}_{B}(\hat{{\gamma}})=\sum_{j}{u}_{\text{\tiny$\bullet$}j}(\hat{{\gamma}}){u}_{\text{\tiny$\bullet$}j}^{\mathrm{\scriptscriptstyle T}}(\hat{{\gamma}}),\quad\hat{V}_{A\cap B}(\hat{{\gamma}})=\sum_{ij}Z_{ij}{u}_{ij}(\hat{{\gamma}}){u}_{ij}^{\mathrm{\scriptscriptstyle T}}(\hat{{\gamma}}),$
where ${u}_{i\text{\tiny$\bullet$}}({\gamma})=\sum_{j}Z_{ij}{u}_{ij}({\gamma})$ and ${u}_{\text{\tiny$\bullet$}j}({\gamma})=\sum_{i}Z_{ij}{u}_{ij}({\gamma}).$ Estimators of the form $\hat{V}_{\text{all}}(\hat{{\gamma}})$ are used in statistical modeling of data clustered within multiple levels in medical applications (Miglioretti & Heagerty, 2004) and in economics (Cameron et al., 2011) where they are known as two-way cluster-robust sandwich estimators.
Finally, we approximate the variance of $\hat{{\beta}}$ by plugging in the estimates $\hat{\sigma}_{A}^{2}$ and $\hat{\sigma}_{B}^{2}$,
$\hat{\mathrm{var}}(\hat{{\beta}})=(1+\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})\,\hat{\mathrm{var}}(\hat{{\gamma}})=(1+\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})\,{J}_{\text{all}}^{-1}(\hat{{\gamma}})\hat{V}_{\text{all}}(\hat{{\gamma}}){J}_{\text{all}}^{-1}(\hat{{\gamma}}).$ (13)
@@ 3 Simulations
In this section, we simulate from the probit model with crossed random effects (1) and compare the performance of the all-row-column estimator with the traditional estimator obtained by maximizing the first-order Laplace approximation of the likelihood. A further method considered is an ‘infeasible oracle estimator’ that uses the unknown true values of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ to estimate the regression parameters as $\hat{{\beta}}_{\mathrm{oracle}}=\hat{{\gamma}}(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}.$ The all-row-column method instead corrects $\hat{{\gamma}}$ using estimates of the variance components. All methods were implemented in the R language. The package TMB (Kristensen et al., 2016) was used for the Laplace approximation, with the nlminb optimization function employed for its maximization. We did not use the popular glmer function from the R package lme4 (Bates et al., 2015) because the current version of TMB is substantially more computationally efficient and thus allowed us to compare our method to the first-order Laplace approximation at larger dimensions than it would otherwise be possible with glmer. The row- and column-wise likelihoods were coded in C++ and integrated in R with Rcpp (Eddelbuettel, 2013), and optimized by Brent’s method as implemented in the optimise function.
@@ 3.1 Simulation settings
We considered eight different settings defined by combining three binary factors. The first factor is whether the simulation is balanced (equal numbers of rows and columns) or imbalanced (with very unequal numbers) like we typically see in applications. The second factor is whether the regression model is null apart from a nonzero intercept or has nonzero regression coefficients. The third factor is whether the random effect variances are set at a high level or at a low level. Given $R$ and $C$, the set $\mathcal{S}$ is obtained by independent and identically distributed Bernoulli sampling with probability $1.27\times N/(RC)$. The value $1.27$ is the largest one for which Ghosh et al. (2022a) could prove that backfitting takes $O(1)$ iterations. This sampling makes the attained value of $N$ random but with a very small coefficient of variation.
@@
We denote by $R=N^{\rho}$ and $C=N^{\kappa}$ the number of rows and columns in the data, as a power of the total sample size $N$. The two levels of the balance factor are thereafter termed ‘balanced’ with $\rho=\kappa=0.56$ and ‘imbalanced’ with $\rho=0.88$ and $\kappa=0.53$. The balanced case was used in Ghosh et al. (2022b), and the imbalanced case is similar to the Stitch Fix data in §4. Because $\rho+\kappa>1$, the fraction of possible observations in the data is $N/(RC)=N^{1-\rho-\kappa}\to 0$ as $N\to\infty$ providing asymptotic sparsity in both cases. While the first choice has $R/C$ constant, the second choice has $R/C\to\infty$ with $N$. We believe that this asymptote is a better description of our motivating problems than either a setting with $C$ fixed as $R\to\infty$ or the setting common in random matrix theory (Edelman & Rao, 2005) where $R$ and $C$ diverge with $R/C$ approaching a constant value. Notice that when $\rho=0.88$ and $\kappa=0.53$ the condition $\max(\rho+2\kappa,2\rho+\kappa)<2$ invoked for the estimator in Ghosh et al. (2022b) does not hold.
@@
We considered seven predictors generated from a multivariate zero-mean normal distribution with a covariance matrix $\Sigma$ corresponding to an autocorrelation process of order one, so that the entry $(k,l)$ of $\Sigma$ is $\phi^{|k-l|}$. We set $\phi=0.5$ in all the simulations. We always used the intercept $\beta_{0}=-1.2$ because in our applications $\mathrm{pr}(Y=1)<1/2$ is typical. For the predictor coefficients we considered two choices termed ‘null’ with $\beta_{\ell}=0$ and ‘linear’ with $\beta_{\ell}=-1.2+0.3\ell$ $(\ell=1,\dots,7).$ The first setting is a null one where ${x}$ is not predictive at all, while the second setting has modestly important nonzero predictors whose values are in linear progression.
@@
The two choices for the variance component parameters are termed ‘high variance’ with $\sigma_{A}=1$ and $\sigma_{B}=1$ and ‘low variance’ with $\sigma_{A}=0.5$ and $\sigma_{B}=0.2$. We chose the first setting to include variances higher than what is typically observed in applications. The second setting is closer to what we have seen in data such as that in §4. We represent the eight settings with mnemonics as shown in Table 1. For example, Imb-Nul-Hi  means row-column imbalance ($\rho=0.88,\kappa=0.53$), all predictor coefficients are zero and the main effect variances are large ($\sigma_{A}=1,\sigma_{B}=1$).
@@
As suggested by a referee, we experimented with different values for the number of quadrature nodes $k$ to approximate the univariate integrals of the row- and column-likelihoods; see also the recent work Bilodeau et al. (2024). The value $k=1,$ which corresponds to the Laplace approximation, produced estimates of $\sigma_{A}$ and $\sigma_{B}$ affected by substantial downward bias even at large $N$. The bias disappeared by increasing $k$. With $k=5,$ we obtained results not affected by bias and essentially indistinguishable from those with $k=25$. Therefore, the results discussed in the rest of this section were calculated with $k=5$ nodes.